-6-(a+9)+7=3a+2(4a-1)
Answer by waynest(281)

-6 -(a + 9) + 7 = 3a + 2(4a -1)
-6 -a -9 + 7 = 3a + 8a -2
-15 + 7 -a = 11a - 2
-8 - a + a + 2 = 11a + a - 2 + 2
-6 = 12a
-6/12 = 12a/12
-1/2 = a
check:
-6 - (-1/2 + 9) + 7 = 3(-1/2) + 2(4(-1/2)- 1)
-6 - (8.5) + 7 = -3/2 - 2(-3)
-6 - 8.5 + 7 = -3/2 - 6
-7.5 = -7.5